OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <8, 8, -6>; v = <3, -8, -3>; w = <-4, -2, -6>

terenzreignz (terenzreignz):

Dot products? :D They're the easier product. You just multiply corresponding components, and then add them. For example, r⋅v. r⋅v = <8,8,-6> ⋅ <3,-8,-3> = 8⋅3 + 8⋅-8 + -6⋅-3 = 24 - 64 + 18 = -22 :) Try v⋅w now :D

OpenStudy (anonymous):

= 3⋅-4 + -8⋅-2 + -3⋅-6 = -12 +16 + 18 = 22 ??